OpenStudy (iesmee):

Solve system by elimination 6x+2y=20 8x+10y=12

OpenStudy (marigirl):

lets eliminate the y variable by multiplying the first equation by 5 \[30x+10y=100\] \[8x+10y=12\]

OpenStudy (marigirl):

now lets subtract the first equation by the second

OpenStudy (marigirl):

30x+10y=100 8x+10y=12 - 22x = 88 x =?

OpenStudy (iesmee):

i subtract 88 by what..?

OpenStudy (marigirl):

we are working our way down and subtracting... 30x+10y=100 8x+10y=12 - 30x-8x=22x + 10y-10y =0 (eliminated) = 100-12=88

OpenStudy (marigirl):

therefore what is left after subtracting one equation from another is 22x=88 x=4

OpenStudy (marigirl):

now sub in x=4 into any equation to find the y value

OpenStudy (iesmee):

ah subtracting by one another

OpenStudy (marigirl):

what is ur y value

OpenStudy (iesmee):

it would be y= 16?

OpenStudy (marigirl):

lets substitute x=4 into any of the two original equations. 6x+2y=20 6(4)+2y=20 now solve for y

OpenStudy (iesmee):

alright

OpenStudy (iesmee):

6 x 4 = 24 plus 2 =26?

OpenStudy (marigirl):

24+2y=20 2y=20-24 2y=-4 y=-2

OpenStudy (iesmee):

you subtract you don't keep adding oh wow

OpenStudy (marigirl):

So both these equations 6x+2y=20 8x+10y=12 cross each other at (4, -2)

OpenStudy (marigirl):

we substituted the x value we found from using the method of elimination

OpenStudy (marigirl):

hmmm i can see how this might get confusing. happy to help if u have questions

OpenStudy (iesmee):

so they must always cross, yes?

OpenStudy (iesmee):

thanks a lot

OpenStudy (marigirl):

i am not sure what u mean by cross. But the story is 6x+2y=20 and 8x+10y=12 are two straight lines and they cross each other at (4, -2). We can find out which points they both have in common by wither elimination or substitution. with elimination we basicly got rid of one of the variables and solved to see what value we get. then substituted that value back into the original equation to find the other point... hope that makes sense.